7) The air in a house is at 20°C and 50 percent relative humidity. Now the air is cooled at constant pressure. The temperature at which the moisture in the air will start condensing 1S (a) 8.7°C (b) 11.3°C (c) 13.8°C (d) 9.3°C (e) 10.0°C On the psychrometric chart, a cooling and dehumidification process 8)a line that is appears as a line that is (a) horizontal to the left, (b) vertical downward, (c) diagonal upwards to the right (NE direction) (d) diagonal upwards to the left (NW direction) (e) diagonal downwards to the left (SW direction) On the psychrometric chart, a simple heating process (no humidification or dehumidification) appears as a line that is (a) horizontal to the right, (b) horizontal to the left, (e) vertical upward, (d) diagonal upwards to the right (NE direction) (e) diagonal downwards to the right (SE direction)
Answer #1

1. The temperature (dew point temperature Td) at which the moisture in the air will start condensing is 9.3

Explanation: there are two way to find dew point temperature Td

i) put the values in the formula given

lnRH 100 b+T RH aT 7 aln

where:

a = 17.625

b = 243.04

T is in ºC

Td is in ºC

RH is in %

T_{d} = \frac{243.04\left ( ln(\frac{50}{100}) \right ) + \frac{17.625*20}{243.06+20}}{17.625-\left ( ln(\frac{50}{100}) \right ) - \frac{17.625*20}{243.06+20}}

Td = 9.20C

ii) alternative method

\phi = RH / 100% = 50/100 = 0.5

where

\phi = Relative Humidity

RH = Relative Humidity in %

saturated water vapour pressure mbar temperature in C C123 Magnus coefficient from table mbar Pa partial water vapour pressur

2. in psychrometric chart cooling and dehumidification process appears as a line that is diagonal downwards to the left (SW direction)

Explanation: When moist air drops below its dew-point by having contact with a cold surface, some amount of the water vapor in the air condenses and allows to convert the air stream as liquid, due to this both the temperature and humidity ratio of air decreases. the actual process path depends upon the surface temperature, cold surface type, and flow conditions, however for simplification of the process,it is assumed to be a straight line. the process can be shown by line O-C

3. in psychrometric chart, a simple heating process (no humidification or dehumidification) appears as a line that is horizontal to the right.

Explanation: when there is a simple heating process, the moisture content of air remains constant (i.e. no humidification or dehumidification) and its temperature increases as it flows over a heating coil is also known as Sensible heating process. it can be shown as process (O-B)